# SPICE[Uncaptioned image]: An Automated SWE-Bench Labeling Pipeline for Issue Clarity, Test Coverage, and Effort Estimation

### v4

@@ V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-V?
Figure 4: Issue labeling (left) and test labeling (right) rationale word count comparison for Human Labelers from SWE-V and SPICE labelers.
RQ3: What are the cost considerations associated with SPICE?
@@
Motivation. As FM-based labeling becomes viable for large-scale SE datasets (e.g., the 21.5k-instance SWE-bench and its multi-thousand-instance variants), practitioners face a crucial quality vs cost dilemma. While SPICE unlocks high-throughput annotation, each inference incurs real monetary and time expenses. Quantifying these expenses across model choices, token volumes, and latency profiles enables informed decisions about which FM configurations balance accuracy with budget and turnaround requirements. This analysis provides actionable guidance to users of our tool for efficiently labeling SWE-like data at scale, while avoiding unexpected financial or time overruns.
@@
Approach. To frame our cost-and-time analysis for this research question, we decompose the evaluation into three complementary components. First, we establish a human-annotation baseline by estimating the per-instance effort and market rates for senior engineers. Next, we capture SPICE’s own operational parameters, including token usage and inference latencies, using our logging infrastructure. Finally, we translate those metrics into API pricing estimates based on public rate cards. Each part is explained in detail below:
– Baseline cost estimation. Here, we estimate the cost of labeling a high quality dataset, such as SWE-V. To estimate the cost of labeling, we obtain the average salaries of a senior software engineer, along with the average time taken to label. Senior software engineers on Upwork bill between \$70 and \$150 per hour [36] and many profiles list rates of \$45 to \$75 per hour [37]. Based on employee-submitted data and internal research, employment websites such as Indeeed and Glassdoor report average rates of approximately \$75/hr [38] and \$81/hr [39] respectively for Upwork senior engineers.
To estimate the labeling time per instance, we reference OpenAI’s reported duration of "roughly 20 minutes per sample" [1]. This also aligns with our own experience labeling SWE-bench-like instances (see Section III).
Let’s assume we need to label ($n$) instances, each by ($m$) annotators at an average rate of ($c$) \$/hr, requiring ($t$) min per instance, with uncertainty factors $\eta_{c},\eta_{t}\in[-10\%,10\%]$. Then:
$C\;=\;n\,m\,(c+\eta_{c})\,\frac{t+\eta_{t}}{60}.$
For n=1,000,m=4,c=\$75,t=20minn=1,000,m=4,c=\\$75,t=20min, the central cost estimate is \$100,000\\$100,000 ($\eta_{c}=\eta_{t}=0$), the minimum cost $C_{\min}$ is \$81,000\\$81,000 ($\eta_{c}=-10\%,\eta_{t}=-10\%$) and the maximum cost $C_{\max}$ is \$121,000\\$121,000 ($\eta_{c}=10,\eta_{t}=10$). In other words, our projected cost for labeling 1,000 instances ranges from \$81k to \$121k, with a central value of \$100k. The estimate considers $n=4$, given that OpenAI’s-SWV [1] involved three labelers, and one moderator. While the moderator can be one of the labelers, they would still have invested cognition and time to reason through the other two labeler’s rationales to decide a final label, and note down her justification for her final choice. Hence, we assume that the average time of 20 minutes also applies for the moderator task as well.
@@
– Parameter Estimation for SPICE. Our SPICE tool leverages litellm [40] as a dependency and avails all the benefits of quick logging. We obtain the issue logs, along with the Aider logs for test logging, and obtain both input and output tokens (the distinction is important because of different pricing), and the time taken for inference calls along with the internal tool working (e.g., querying the relevant files by Aider). For all estimates, like input and output tokens, we used SPICEG-4o-m for ICA and SPICEDS-R for TCA. However, while estimating time, we did not use Deepseek API, since Deepseek’s inference can fluctuate based on server load [41, 42, 43, 44]; preventing us from getting an accurate estimate. Hence, we used the time logs from queries to OpenAI API using SPICEG-4o-m for both ICA and TCA for estimating time.
@@
– Estimation of API pricing. We obtain input and output cost values from public pricing pages [45, 46]. While server-side caching may reduce effective cost, it cannot be computed/estimated at the client side, and not all FM providers provide this, we exclude it from our estimates, and report the worst case scenarios.
@@
Results. SPICE sends a notably high amount (median 75,000) of input tokens (i.e., prompt tokens) for TCA and receives a median of 4,700 output tokens (i.e., model response tokens). Figure 5 illustrates the token distribution for issue and test labeling. The high number of output tokens for TCA arises because SPICE leverages additional contextual information from multiple related files associated with each patch. In contrast, ICA labeling requires significantly fewer tokens, with a median of only 3,162 output tokens per instance and a median of 475 input tokens per instance. This is also reflected in the time taken for generating issue labels compared to test labels. Labeling one issue takes a median time of 10 second, whereas labeling one test is approximately 8 times slower at 83 seconds. Naturally, the actual time is dependent on server load, network bandwidth, etc., and can vary.
Overall,the total time taken by SPICE is approximately 53 times lower than that taken by humans: we estimate a total of 1,333 engineer-hours to label 1K instances (i.e., 20 min per instance * 4-expert panel * 1,000 instances), as opposed to 25 hours (i.e., approximately 90 sec per instance * 1000 instances) by SPICE.
@@
Table II provides the estimated labeling costs for 1,000 instances based on median token usage, should these models be used for both issue and test labeling. These costs combine total input and output tokens with their respective pricing structures. As per our default setting (SPICEG-4o-m for ICA and SPICEDS-R for TCA), labeling 1,000 instances using SPICE costs approximately \$5.1 (\$0.11 for issues and \$4.99 for tests). In contrast, our baseline cost using four senior software engineers is \$100,000. This translates to a remarkable cost reduction factor of approximately 19,600 times, highlighting SPICE’s extraordinary cost efficiency.
@@
Offline quantized models, such as Qwen-14B achieves approximately 75% accuracy in ICA. However, their accuracy drops significantly below 45% for test adequacy assessments. Consequently, these locally hosted models can be best leveraged as cost-effective preliminary tools for simpler ICA labeling, rather than the TCA labeling.
Additional cost savings can be achieved by modifying SPICE’s default parameters, such as the number of SPICE labelers, assigning different models to different SPICE labelers, or limiting the size of the code context retrieved by Aider (which SPICE conveniently supports via configurable settings). The users of our tool may wish to evaluate the cost/performance trade-offs by tinkering around with these parameters.
VI Related Work
@@
SWE-bench Limitations and Variants. While SWE-V (a subset of SWE-bench) largely superseded SWE-bench, it still has clear limitations. The dataset is relatively small (500 instances), all 12 projects are written in Python, and empirical evidence suggests that newer models were trained with SWE-bench data and have memorized issue-fix pairs [59] (SPICE can be used to create fresh datasets and mitigate the risks of benchmark data leakage during pre-training). Other research studies have recently flagged additional data quality issues in SWE-V [18]. Importantly, SWE-V patches do not cover all corner cases, and incorrect patches can still sneak through [60].
As a response to the aforementioned limitations, several SWE-bench-like datasets have been proposed. For example, in terms of broadening the scope of SWE-V, ByteDance’s SWE-bench-Multilingual [3] includes 300 tasks in nine languages and SWE-bench Multimodal [4] include approximately 600 GUI-based JavaScript tasks with screenshots. A non-exhaustive list of SWE-V variants include: SWE-Bench+ [18], SWE-Lancer [61], SWE-Gym [2], SWE-bench-Live [59], and SWE-bench Lite.
@@ VII Threats to Validity
In addition, although SPICE can estimate task difficulty, we do not evaluate it in our research questions, because our primary objective is selecting solvable issues rather than measuring annotator effort. Moreover, testing is costly. Each additional experiment adds non-trivial monetary expense.
However, in the working of our tool, since difficulty-estimation procedure is exactly the same as TCA’s except for a modified prompt, we assume that our RQ2 and RQ3 results will be similar to that of TCA’s. Specifically, SPICE difficulty rationales in RQ2 should exhibit the similar (7x) increase in length, and that token usage and cost projections in RQ3 should be comparable. Although the evaluation of our tool’s full capabilities remains a threat, it does not diminish our tool’s value in filtering of “unsolvable” issues. Moreover, difficulty labels are provided in our SPICE-bench release, along with the evaluation prompt to support future work.

### v5

@@ V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-L?
Figure 5: Issue labeling (left) and test labeling (right) rationale word count comparison for Human Labelers from SWE-L and SPICE labelers.
V-C RQ3: What are the cost considerations associated with SPICE?
@@
Motivation. As FM-based labeling becomes viable for large-scale SE datasets (e.g., the 21.5k-instance SWE-bench and its multi-thousand-instance variants), practitioners face a crucial quality vs cost dilemma. While SPICE unlocks high-throughput annotation, each inference incurs real monetary and time expenses. Quantifying these expenses across model choices, token volumes, and latency profiles enables informed decisions about which FM configurations balance accuracy with budget and turnaround requirements. This analysis provides actionable guidance to users of our tool for efficiently labeling SWE-bench-like data at scale, while avoiding unexpected financial or time overruns.
@@
Approach. To frame our cost-and-time analysis for this research question, we decompose the evaluation into two complementary components. First, we capture SPICE’s own operational parameters, including token usage and inference latencies, using our logging infrastructure. Then, we translate those metrics into API pricing estimates based on public rate cards. Both parts are explained in detail below:
@@
– Parameter Estimation for SPICE. Our SPICE tool leverages litellm [36] as a dependency and avails all the benefits of quick logging. We obtain the issue logs, along with the Aider logs for test logging, and obtain both input and output tokens (the distinction is important because of different pricing), and the time taken for inference calls along with the internal tool working (e.g., querying the relevant files by Aider). For all estimates, like input and output tokens, we used SPICEG-4o-m for ICA and SPICEDS-R for TCA (i.e., we use the same 110 instances from RQ1 and RQ2). While estimating time, we did not use Deepseek API, since Deepseek’s inference can fluctuate based on server load [37, 38, 39, 40]; preventing us from getting an accurate estimate. Hence, we used the time logs from queries to OpenAI API using SPICEG-4o-m for both ICA and TCA for estimating time.
@@
– Estimation of API pricing. We obtain input and output cost values from public pricing pages [41, 42]. While server‑side caching may reduce effective cost, it cannot be computed/estimated at the client side, and not all FM providers provide this, we exclude it from our estimates, and report the worst case scenarios.
@@
Results. SPICE sends a notably high amount (median 75,000) of input tokens (i.e., prompt tokens) for TCA and receives a median of 4,700 output tokens (i.e., model response tokens). Figure 6 illustrates the token distribution for issue and test labeling. The high number of output tokens for TCA arises because SPICE leverages additional contextual information from multiple related files associated with each patch. In contrast, ICA labeling requires significantly fewer tokens, with a median of only 3,162 output tokens per instance and a median of 475 input tokens per instance. This is also reflected in the time taken for generating issue labels compared to test labels. Labeling one issue takes a median time of 10 seconds, whereas labeling one test is approximately 8 times slower at 83 seconds. Naturally, the actual time is dependent on server load, network bandwidth, etc., and can vary.
@@
Table II provides the estimated labeling costs for 1,000 instances based on median token usage, should these models be used for both issue and test labeling. These costs combine total input and output tokens with their respective pricing structures. As per our default setting (SPICEG-4o-m for ICA and SPICEDS-R for TCA), labeling 1,000 instances using SPICE costs approximately \$5.1 (\$0.11 for issues and \$4.99 for tests).
@@
Additional measures can be taken for reducing costs. From our experiments, we observe that offline quantized models performed reasonably well in ICA. For instance, SPICEQ-14 achieves approximately 75% accuracy in ICA, which is only 12 percentage points below the top-performing model (Table I). Additional cost savings can be achieved by modifying SPICE’s default parameters, such as the number of SPICE labelers, assigning different models to different SPICE labelers, or limiting the size of the code context retrieved by Aider (which SPICE conveniently supports via configurable settings). The users of our tool may wish to evaluate the cost/performance trade-offs by tinkering around with these parameters.
VI The SPICE[Uncaptioned image]-Bench Dataset
A core contribution of this paper is the release of SPICE-Bench. SPICE-Bench comprises 6,802 SPICE-labeled instances (code-repair tasks) collected from 291 repositories sampled from the SWE-Gym dataset[2]. This number of repositories is much larger than that of SWE-V (12), making SPICE-Bench significantly more diverse and heterogeneous.
58.5% (3,981) of the instances in SPICE-Bench contain a well-specified issue, while 49.2% (3,342) contain adequate test coverage. 31.8% (2,164) of instances contain both a well-specified issue and adequate test coverage. This set of 2,164 “good” instances is 4.3x the size of SWE-V (500).
Analogously to SWE-L, SPICE-Bench also includes labels for difficulty (0 indicates > 15 mins of effort, 1 indicates 15 mins-1hr, 2 indicates 1-4 hrs, and 3 indicates 4+ hrs.). Out of the 2,164 good instances, 7.8% (169) scored 0, 27.8% (602) scored 1, 60.9% (1,218) scored 2, and 3.5% (75) scored 3 for difficulty.
More information about SPICE[Uncaptioned image]-Bench is available on Hugging Face*444*The URL will be made available after acceptance..
VII Related Work
@@
SWE-bench Limitations and Variants. While SWE-V largely superseded SWE-bench, it still has clear limitations. The dataset is relatively small (500 instances), all 12 projects are written in Python, and empirical evidence suggests that newer models were trained with SWE-bench data and have memorized issue-fix pairs [55] (SPICE can be used to create fresh datasets and mitigate the risks of benchmark data leakage during pre-training). Other research studies have recently flagged additional data quality issues in SWE-V [18]. For instance, Wang et al. [56] indicate that SWE-V patches do not cover all corner cases and incorrect patches can still sneak through. The authors call for more robust evaluation practices, including running all available tests, filtering incorrect patches, and strengthening benchmarks with differentiating tests generated by their proposed tool (PatchDiff) to build a sustainable patch validation ecosystem.
As a response to the aforementioned limitations, several SWE-bench-like datasets have been proposed. For example, in terms of broadening the scope of SWE-V, ByteDance’s SWE‑bench-Multilingual [3] includes 300 tasks in nine languages and SWE‑bench Multimodal [4] include approximately 600 GUI-based JavaScript tasks with screenshots. A non-exhaustive list of SWE-V variants include: SWE-Bench+ [18], SWE‑Lancer [57], SWE‑Gym [2], SWE-bench-Live [55], and SWE-bench Lite.
@@ VIII Threats to Validity
In addition, although SPICE can estimate task difficulty, we do not evaluate it in our research questions, because our primary objective is selecting solvable issues rather than measuring annotator effort. Moreover, testing is costly. Each additional experiment adds non‐trivial monetary expense.
However, in the working of our tool, since difficulty‐estimation procedure is exactly the same as TCA’s except for a modified prompt, we assume that our RQ2 and RQ3 results will be similar to that of TCA’s. Specifically, SPICE difficulty rationales in RQ2 should exhibit the similar (7x) increase in length, and that token usage and cost projections in RQ3 should be comparable. Although the evaluation of our tool’s full capabilities remains a threat, it does not diminish our tool’s value in filtering of “unsolvable” issues. Moreover, difficulty labels are provided in our SPICE-bench release, along with the evaluation prompt to support future work.